# Machine Learning Detection of Lithium Plating in Lithium-ion Cells: A Gaussian Process Approach

### v1

@@ I INTRODUCTION
Recent studies have reported new voltage features that emerge during charging. In particular, differentiating the accumulated charge $Q$ (the integral of current) with respect to voltage $V$ yields a $dQ/dV$ curve that exhibits a distinct peak above a threshold voltage, which has been directly linked to the onset of lithium plating [12, 11]. Operando experiments using novel capillary cell setups have confirmed this finding by directly visualizing the growth of metallic Li deposits [12, 13]. This discovery opens a promising pathway toward real-time detection and control of lithium plating: in principle, one could compute $dQ/dV$ from routine current and voltage measurements and terminate charging when the peak is observed. However, practical implementation poses significant challenges. The peak in the $dQ/dV$ curve reflects the flattening of the voltage trajectory, since mathematically $dQ/dV$ is the inverse of the time derivative of voltage under constant-current charging. As is well known in control theory, numerical differentiation of signals (e.g., via finite differencing) significantly amplifies sensor noise, leading to large variance in the estimated derivative. While laboratory demonstrations can mitigate this issue with high-precision instrumentation and extensive post-processing (e.g., filtering), robust and accurate detection of $dQ/dV$ peaks based on near-zero voltage derivatives remains extremely difficult in practical settings.
@@ II Mechanism and Lithium Plating and Electrochemical Features
Lithium plating occurs when, during charging, lithium ions are reduced and deposited as metallic lithium on the graphite anode rather than intercalating into its structure. This parasitic process is favored under conditions of high current or low temperature, where polarization drives the anode potential sufficiently negative relative to Li/Li+. Once initiated, plating accelerates rapidly, consuming cyclable lithium and producing deposits that may evolve into dendrites capable of piercing the separator and triggering internal short circuits. As a result, lithium plating represents a uniquely fast and catastrophic degradation pathway compared with gradual mechanisms such as solid-electrolyte interphase (SEI) growth.
Differential capacity analysis, or $dQ/dV$ analysis, has long been a standard tool for diagnosing battery degradation. Under low-rate, near-equilibrium charging conditions, the VVQQ curve reflects the thermodynamics of electrode reactions, and its derivative produces a spectrum of peaks corresponding to distinct electrochemical processes such as graphite staging transitions or cathode phase changes. Over aging, systematic peak shifts and broadening provide sensitive markers for gradual degradation mechanisms, including loss of active material (LAM) and loss of lithium inventory (LLI).
Lithium plating, however, differs fundamentally from these slow mechanisms. It progresses rapidly and can cause severe capacity loss within only a few charge–discharge cycles. Failure to detect plating early not only shortens cell lifetime but also introduces serious safety risks.
Recent experimental studies have revealed that lithium plating produces a distinctive new peak in the incremental capacity spectrum, typically appearing above $4.0~\mathrm{V}$ across a wide range of cell chemistries. This feature arises from the abrupt change in current partitioning once the anode potential crosses the lithium deposition threshold. Consequently, the presence of a $dQ/dV$ peak above $4.0~\mathrm{V}$ has emerged as a strong electrochemical signature for the onset of plating.
Despite its promise, applying $dQ/dV$ analysis for plating detection under practical fast-charging conditions remains challenging. At high C-rates, overpotentials and IR drop distort the voltage curve, shifting and broadening differential features. More critically, numerical differentiation of noisy voltage–current measurements amplifies sensor noise, producing fluctuations that obscure the plating peak. Traditional filtering methods such as Savitzky-Golay [16], can mitigate variance but introduce their own biases, making robust peak detection under realistic charging conditions an open challenge.
@@ III Gaussian Process-based Lithium Plating Detection Framework
Gaussian Processes (GPs) offer a principled alternative: they model the underlying charge–voltage map as a smooth stochastic function with calibrated uncertainty, and enable noise-robust derivative inference via closed-form conditioning rather than unstable differencing [14, 15, 18]. Concretely, GPs (i) capture nonlinear yet smooth electrochemical responses; (ii) provide credible intervals that quantify confidence in the presence and location of plating-related peaks; and (iii) admit scalable variants (e.g., state-space, sparse, or variational GPs) suitable for online deployment in battery management systems (BMS) for both EVs and spacecraft [19, 20, 21].
@@ III-B Inference of $dQ/dV$
Theorem 1 (Derivative GP and Joint Gaussianity).
@@ IV-A Testing Setup
To validate the proposed lithium plating detection scheme, we conducted constant-current (CC) charge/discharge aging cycles under multiple operating conditions. Charging was performed at C-rates of 1C, 0.8C, 0.6C, 0.4C, and 0.2C, across ambient temperatures ranging from $40\,^{\circ}\text{C}$ down to $0\,^{\circ}\text{C}$ (inclusive). Each charging cycle was followed by discharging at 1C down to the manufacturer-specified lower cutoff voltage $V_{\min}=2.75~\text{V}$, with an upper cutoff voltage $V_{\max}=4.2~\text{V}$. A rest period of $5~\text{min}$ was inserted at the end of each cycle, and at least 10 cycles were repeated under each condition. The experiments were performed using an Arbin LBT 21084 battery cycler for charge/discharge control and data acquisition, together with an ESPEC environmental chamber for temperature regulation. The tested cell was a 45 mAh LIR 2032 Li-ion coin cell. To track degradation, periodic Reference Performance Tests (RPTs) were conducted, including 1/10 C constant-current capacity tests and hybrid pulse power characterization (HPPC) to evaluate internal resistance. A typical testing sequence is shown in Fig. 1.
@@ IV-B Results and Discussion
Figure 2: GP-based plating detection across charge rates and temperatures. (a) V(t) during CC charge for representative cycles (legend shows C-rate and temperature). (b) capacity–voltage trajectories Q(V). (c) GP-derived incremental capacity dQ/dV with inset figures (i), (ii) and (iii) showing the magnified 95% credible bands at the start of the cycle with bounds as % of peak magnitude of the no-plating and plating clusters, respectively.
@@
Using voltage–current data collected during CC charging, we trained GP models on the measured $Q(V)$ trajectory and inferred $dQ/dV$ directly from the derivative GP posterior. This probabilistic differentiator yields smooth incremental-capacity curves with calibrated uncertainty, enabling robust peak finding under realistic sensor noise. Fig. 2 illustrates the full GP-based analysis workflow for several charge-rate/temperature combinations. The first subplot shows $V(t)$ during CC charge, where increasing C-rate and decreasing temperature shorten charge time due to increased polarization. The second subplot maps the data to $Q(V)$ and overlays the GP posterior. The third subplot presents the GP-derived $dQ/dV$ (mean and $\pm$95% credible interval), revealing transient differential peaks that are otherwise obscured by finite-difference noise.
@@
Table I summarizes all the findings. In plating cases, charge throughput and capacity loss (as a %/cycle) are lower by about an order of magnitude compared with non-plating cases. Our GP-based method successfully identified peaks above $4.0~\mathrm{V}$ in all cases where plating occurred, while correctly reporting no peaks under non-plating conditions. Taken together, (i) the presence of $dQ/dV$ peaks above $4.0~\mathrm{V}$, (ii) the observed reduction in charge throughput over aging cycles, and (iii) the increased capacity fade confirmed by RPT measurements validate the accurate detection capability of our method.
Finally, it is important to note the role of GP uncertainty quantification. Across all plating cases, the credible intervals around the detected peaks were exceptionally tight. On average, the bounds were within 0.3% of the peak magnitude, indicating very high confidence in the detection. This level of certainty is unattainable with conventional methods, which provide only smoothed (point) estimates without any measure of confidence.
@@ V Conclusions
These differential signatures align closely with independent performance metrics. Plating-positive cases exhibit faster declines in delivered charge throughput per CC cycle as well as accelerated capacity fade confirmed by periodic RPTs, with the coldest and highest-rate cases degrading most severely. In contrast, baseline and non-plating cases show stable charge throughput and much slower capacity fade.
@@ ACKNOWLEDGMENT
We appreciate the funding support from the the AFWERX Proteus Space STTR (Grant No.XXXXX) and the NSF CAREER Program (Grant No.2046292).

### v2

@@ I INTRODUCTION
Recent studies have reported new voltage features that emerge during charging [12, 13]. In particular, differentiating the accumulated charge $Q$ (the integral of current) with respect to voltage $V$ yields a $dQ/dV$ curve that exhibits a distinct peak above a threshold voltage, which has been directly linked to the onset of lithium plating. Operando experiments using novel capillary cell setups have confirmed this finding by directly visualizing the growth of metallic Li deposits [12]. This discovery opens up a promising pathway toward real-time detection and control of lithium plating: in principle, one could compute $dQ/dV$ from routine current and voltage measurements and terminate charging when the peak is observed. However, practical implementation poses significant challenges. The peak in the $dQ/dV$ curve reflects the flattening of the voltage trajectory, since mathematically $dQ/dV$ is the inverse of the time derivative of voltage under constant-current charging. As is well known in control theory, numerical differentiation of signals (e.g., via finite differencing) significantly amplifies sensor noise, leading to large variance in the estimated derivative. While laboratory demonstrations can mitigate this issue with high-precision instrumentation and extensive post-processing (e.g., filtering), robust and accurate detection of $dQ/dV$ peaks based on near-zero voltage derivatives remains extremely difficult in practical settings.
@@ II Mechanism and Lithium Plating and Electrochemical Features
Lithium plating occurs when, during charging, lithium ions are reduced and deposited as metallic lithium on the graphite anode rather than intercalating into its structure [2]. This parasitic process is favored under conditions of high current or low temperature, where polarization drives the anode potential sufficiently negative relative to Li/Li+. Once initiated, plating accelerates rapidly, consuming cyclable lithium and producing deposits that may evolve into dendrites capable of piercing the separator and triggering internal short circuits. As a result, lithium plating represents a uniquely fast and catastrophic degradation pathway compared with gradual mechanisms such as solid-electrolyte interphase (SEI) growth.
Incremental capacity analysis, or $dQ/dV$ analysis, has long been a standard tool for diagnosing battery degradation [17]. Under low-rate, near-equilibrium charging conditions, the VVQQ curve reflects the thermodynamics of electrode reactions, and its derivative produces a spectrum of peaks corresponding to distinct electrochemical processes such as graphite staging transitions or cathode phase changes. Over aging, systematic peak shifts and broadening provide sensitive markers for gradual degradation mechanisms, including loss of active material (LAM) and loss of lithium inventory (LLI).
Lithium plating, however, differs fundamentally from these slow mechanisms. It progresses rapidly and can cause severe capacity loss within only a few charge–discharge cycles. Failure to detect plating early not only shortens cell lifetime but also introduces serious safety risks. Recent experimental studies have revealed that lithium plating produces a distinctive new peak in the incremental capacity spectrum, typically appearing above $4.0~\mathrm{V}$ across a wide range of cell chemistries [12, 13]. This feature arises from the abrupt change in current partitioning once the anode potential crosses the lithium deposition threshold. Consequently, the presence of a $dQ/dV$ peak above $4.0~\mathrm{V}$ has emerged as a strong electrochemical signature for the onset of plating.
Despite its promise, applying $dQ/dV$ analysis for plating detection under practical fast-charging conditions remains challenging. At high C-rates, overpotentials and IR drop distort the voltage curve, shifting and broadening differential features. More critically, numerical differentiation of noisy voltage–current measurements amplifies sensor noise, producing fluctuations that obscure the plating peak. Traditional filtering methods such as Savitzky-Golay [16, 17], can mitigate variance but introduce their own biases, making robust peak detection under realistic charging conditions an open challenge.
@@ III Gaussian Process-based Lithium Plating Detection Framework
Gaussian Processes (GPs) offer a principled alternative: they model the underlying charge–voltage map as a smooth stochastic function with calibrated uncertainty, and enable noise-robust derivative inference via closed-form conditioning rather than unstable differencing [14, 15, 19]. Concretely, GPs (i) capture nonlinear yet smooth electrochemical responses; (ii) provide credible intervals that quantify confidence in the presence and location of plating-related peaks; and (iii) admit scalable variants (e.g., state-space, sparse, or variational GPs) suitable for online deployment in battery management systems (BMS) for both EVs and spacecraft [20, 21].
@@ III-B Inference of $dQ/dV$
Theorem 1 (Derivative GP and Joint Gaussianity [14]).
@@ IV-A Testing Setup
To validate the proposed lithium plating detection scheme, we conducted constant-current ($C$) charge/discharge aging cycles under multiple operating conditions. Charging was performed at C-rates of 1C, 0.8C, 0.6C, 0.4C, and 0.2C, across ambient temperatures ranging from $40\,^{\circ}\text{C}$ down to $0\,^{\circ}\text{C}$ (inclusive). Each charging cycle was followed by discharging at 1C down to the manufacturer-specified lower cutoff voltage $V_{\min}=2.75~\text{V}$, with an upper cutoff voltage $V_{\max}=4.2~\text{V}$. A rest period of $5~\text{min}$ was inserted at the end of each cycle, and at least 10 cycles were repeated under each condition. The experiments were performed using an Arbin LBT 21084 battery cycler for charge/discharge control and data acquisition, together with an ESPEC environmental chamber for temperature regulation. The tested cell was a 45 mAh LIR 2032 Li-ion coin cell. To track degradation, periodic Reference Performance Tests (RPTs) were conducted, including 1/10 C constant-current capacity tests and hybrid pulse power characterization (HPPC) to evaluate internal resistance. A typical testing sequence is shown in Fig. 1.
@@ IV-B Results and Discussion
Figure 2: GP-based plating detection across charge rates and temperatures. (a) V(t) during $C$ charge for representative cycles (legend shows C-rate and temperature). (b) capacity–voltage trajectories Q(V). (c) GP-derived incremental capacity dQ/dV with inset figures (i), (ii) and (iii) showing the magnified 95% credible bands at the start of the cycle with bounds as % of peak magnitude of the no-plating and plating clusters, respectively.
@@
Using voltage–current data collected during $C$ charging, we trained GP models on the measured $Q(V)$ trajectory and inferred $dQ/dV$ directly from the derivative GP posterior. This probabilistic differentiator yields smooth incremental-capacity curves with calibrated uncertainty, enabling robust peak finding under realistic sensor noise. Fig. 2 illustrates the full GP-based analysis workflow for several charge-rate/temperature combinations. The first subplot shows $V(t)$ during $C$ charge, where increasing C-rate and decreasing temperature shorten charge time due to increased polarization. The second subplot maps the data to $Q(V)$ and overlays the GP posterior. The third subplot presents the GP-derived $dQ/dV$ (mean and $\pm$95% credible interval), revealing transient differential peaks that are otherwise obscured by finite-difference noise.
@@
Table I summarizes all the findings. In plating cases, charge throughput and capacity loss (as a %/cycle) are higher by about an order of magnitude compared with non-plating cases. Our GP-based method successfully identified peaks above $4.0~\mathrm{V}$ in all cases where plating occurred, while correctly reporting no peaks under non-plating conditions. Taken together, (i) the presence of $dQ/dV$ peaks above $4.0~\mathrm{V}$, (ii) the observed reduction in charge throughput over aging cycles, and (iii) the increased capacity fade confirmed by RPT measurements validate the accurate detection capability of our method.
Finally, it is important to note the role of GP uncertainty quantification. Across all plating cases, the credible intervals around the detected peaks were exceptionally tight as shown in the insets of Fig. 2(c). On average, the bounds were within 0.3% of the peak magnitude, indicating very high confidence in the detection. This level of certainty is unattainable with conventional methods, which provide only smoothed (point) estimates without any measure of confidence.
@@ V Conclusions
These differential signatures align closely with independent performance metrics. Plating-positive cases exhibit faster declines in delivered charge throughput per $C$ cycle as well as accelerated capacity fade confirmed by periodic RPTs, with the coldest and highest-rate cases degrading most severely. In contrast, baseline and non-plating cases show stable charge throughput and much slower capacity fade.
@@ ACKNOWLEDGMENT
We appreciate the funding support from the USAF Proteus Space STTR and the NSF CAREER Program (Grant No.2046292).